40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 206     Entry time: Sat Jul 10 05:31:15 2010
Author: rana 
Type: Summary 
Category: NoiseBudget 
Subject: from the SVN today 

I modified the photo-thermal part of the NB to use the Alnis et. al. number of 25 Hz / uW  (frequency shift per input power noise). I adjusted it for the ratio of the cavity lengths between theirs and ours, and also for the ratio of the Finesses (400000 : 10000).

Tara should now grab this code and replace my guess at the RC RIN with his real measurements and also add the previously measured RC/AC frequency noise to this plot.

The punchline is that the this thermo-refractive noise is significant and we must control the low frequency RIN.

Attachment 1: psl_refcav_sio2_300K.png  71 kB  | Hide | Hide all
psl_refcav_sio2_300K.png
ELOG V3.1.3-